# Constant Modulus Waveforms for IoT-Centric Integrated Sensing and Communications

### v1

@@ Abstract
Integrated sensing and communications (ISAC) is considered a key enabler to support application scenarios such as the Internet-of-Things (IoT) in which both communications and sensing play significant roles. Multi-carrier waveforms, such as orthogonal frequency division multiplexing (OFDM), have been considered as good candidates for ISAC due to their high communications data rate and good time bandwidth property for sensing. Nevertheless, their high peak-to-average-power-ratio (PAPR) values lead to either performance degradation or an increase in system complexity. This can make OFDM unsuitable for IoT applications with insufficient resources in terms of power, system complexity, hardware size or cost. This article provides IoT-centric constant modulus waveform designs that leverage the advantage of unit PAPR and thus are more suitable in resource-limited scenarios. More specifically, several single-carrier frequency and/or phase-modulated waveforms are considered. A comprehensive discussion on their radar sensing and communications performance is conducted based on performance metrics, including the radar ambiguity function, the bandwidth property, the data rate, and the communications receiver complexity.
@@ I-A Standardisation Status
Standardisation plays a crucial role in developing a consistent framework for ISAC in next generation wireless networks, within which both academia and industry can work. Recently, the industry has started laying the vision for the International Telecommunication Union Radiocommunication Sector (ITU-R) International Mobile Telecommunications-2030 (IMT-2030) towards the sixth generation of wireless communications technology (6G). ISAC has emerged as a popular theme in this vision, which led to its adoption as one of the six usage scenarios in ITU-R IMT-2030 [1]. This interest was the culmination of the progress made in researching orthogonal frequency division multiplexing (OFDM)-based radio radar sensing during the last decade, the successful introduction of radio frequency (RF) sensing in the Institute of Electrical and Electronics Engineers (IEEE) 802 standards at both low and high frequencies, and the enhancements made for the positioning framework in the Third Generation Partnership Project (3GPP). This positioning work is part of 5G New Radio (NR) to enable centimetre-level positioning as required by certain vertical industries. Following ISAC adoption in ITU-R IMT-2030, in late 2023, the European Telecommunications Standards Institute (ETSI) launched an industry specification group (ISG) on ISAC with the mission to streamline ISAC pre-standardisation research and pave the way for its specification in 6G standards [2]. In parallel, 3GPP Technical Specification Group (TSG) on Service and System Aspects Working Group 1 (SA1) launched a feasibility study on ISAC use cases in 5G-Advanced [3]. Later in 2024, 3GPP TSG Radio Access Network Working Group 1 (RAN1) started a study item focused on extending the 3GPP baseline channel model in Technical Report (TR) 38.901 to support ISAC and hence pave the way for ISAC specification in both 5G-Advanced and 6G releases. At the recent 3GPP 6G workshop in March 2025, ISAC was confirmed as one of the key themes proposed by the industry for 3GPP to specify in the forthcoming Release 21 (2028-2029) [4]. In addition to the ongoing work of the standards development organisations, extensive pre-standardisation work for important 6G ISAC use cases has also been conducted and published by various important industrial associations. These include the 5G Automotive Association [5], 5G Alliance for Connected Industries and Automation [6] and one6G [7], who have presented various use cases in industrial and/or automated vehicle scenarios, with details of their assumptions, functional and performance requirements.
@@
In the development of the 6G framework, massive and sustainable Internet of Things (IoT) is considered as one of the key usage services for mobile users. ISAC is able to fulfil the sustainability aspect very well since it supports the sharing of spectral and hardware resources between sensing and communication operations. This, in turn, enables emerging IoT applications such as vehicle-to-everything communications, remote sensing, energy management, etc.
@@ I-B Waveform Design
Various approaches have been taken to tackle this joint waveform design problem [8]. Sensing (or radar) centric methods focus on embedding information into traditional radar waveforms, including linear frequency-modulated (LFM), stepped frequency or frequency hopping (FH) multiple-input multiple-output (MIMO) waveforms. Communications-centric designs integrates the radar sensing functionality into existing communications systems. Examples include designing preambles of IEEE 802.11-based frameworks and using multi-carrier waveforms such as OFDM or orthogonal time frequency space (OTFS) for sensing. Waveforms designed from the ground up do not rely on existing radar or communications waveforms. These designs are usually formulated as optimisation problems which consider both communications and sensing performance metrics by allocating resources among temporal, spectral and/or spatial domains. It can be observed from existing methods that the time-frequency resource allocation plays a significant role in tackling the joint waveform design problem. For example, stepped frequency radar waveforms uses a single subcarrier in each time slot, while the frequencies in different time slots are different. Thus, the entire waveform occupies a large bandwidth, leading to a high range estimation resolution. Nevertheless, its multi-carrier counterparts can achieve a much higher data rate while guaranteeing the same large bandwidth. In addition, multi-carrier systems allow for trading off between communications and radar performance by adjusting the portions of subcarriers used for transmitting data and those designed for sensing. These are some of the main reasons why OFDM has been considered an important candidate for ISAC.
Recent work has focused on multi-carrier waveforms such as OFDM as an effective high-rate approach when the required power and computational capability are available. Nevertheless, many 6G systems may not have sufficient resources in terms of power, cost and size to support advanced multi-carrier techniques with the associated circuitry and components. For example, some existing systems such as Bluetooth and long-range radio (LoRA), as well as new applications introduced by 6G, such as internet of radars (IoR), smart wearables and wireless radar sensor networks require low-power, inexpensive devices [9]. Therefore, single carrier modulation and low PAPR waveforms are recommended for such applications that utilise high frequency bands [10]. Given that constant envelope modulation only requires simple, low-cost circuits and operates power amplifiers at the highest efficiency point, it is ideal in low-power systems where energy efficiency and cost outweigh spectrum efficiency. As a result, in this work we focus on constant modulus waveform designs suitable for joint radar and communication.
@@ I-C Contributions
  • We introduce multiple constant modulus frequency and/or phase-modulated waveforms. General discussions on the performance of these waveforms are also conducted.
  • We provide extensive numerical examples to perform a thorough comparison among the constant modulus waveforms based on the introduced performance metrics.
@@ II-A Importance of Low PAPR
Nonetheless, multi-carrier waveforms that suffer from high PAPR have received significant attention as ISAC waveforms due to their high data rate. As a result, different PAPR reduction techniques such as linearisation circuits have also been proposed for ISAC systems to reduce the performance degradation [12]. However, such solutions require additional overheads of circuitry and increased complexity. This, in turn, increases the device cost or degrades the bit-error-rate (BER) performance due to OFDM signal distortion [14], making them unsuitable for low-cost, low-complexity and low-power applications. For example, existing applications such as Bluetooth, LoRA and IoR do not have sufficient resources in terms of power, size and cost to accommodate multi-carrier modulation with complex PAPR reduction techniques. In contrast, constant PAPR waveforms simplify the design process for antennas and RF components, thereby offering advantages such as better performance with simplified RF components, reduced size and weight, and lower cost, making them ideal for applications where energy efficiency and cost outweigh spectrum efficiency.
@@ II-B Radar Sensing Performance
Performance of the radar waveform can be measured in terms of local accuracy and global accuracy. Focusing on the mainlobe of the delay-Doppler plane, we can analyse the local accuracy. It is measured by the delay resolution and the Doppler resolution. The Doppler resolution is determined by the overall time duration of the signal and, it remains constant if the time duration of the waveforms are constant. On the other hand, the radar delay resolution is directly related to the mean square bandwidth of the radar signal under a pulse compression system. The higher the mean square bandwidth, the higher the delay resolution. Therefore, we consider the mean square bandwidth to asses the radar local performance. Focusing on the rest of the delay-Doppler plane we can analyse the global accuracy of radar sensing. It describes the accuracy of target detection in the presence of clutter and this can be measured by false alarm rate and miss-detection probability. When the transmit power remains the same, as is the case for all the waveforms discussed in this paper, the miss-detection probability will be almost the same for a given threshold. On the other hand, the false alarm rate depends on the energy distribution of the received signal. Using an important radar analytical tool known as the ambiguity function (AF), this can be evaluated using the peak sidelobe level (PSL), defined as the ratio of the magnitude of the highest sidelobe peak to the magnitude of the mainlobe peak in the AF. If the PSL is lower, it mitigates the masking effect of nearby targets and reduces the false alarm rate. Therefore, we consider the AF PSL to assess the radar global accuracy.
@@ II-C Communications Performance
From the communication perspective, the performance can be measured in terms of throughput/data rate, reliability and feasibility of implementation. Note that the communication data rate depends on the channel and can vary significantly between different communication channels. Given that the scope of this work is limited to a review of different waveforms, we consider the maximum possible data rate achievable under a perfect channel instead of focusing on data rate under different communication channels. Therefore, we consider the number of bits that can be transmitted in each sub-pulse, irrespective of the channel, to assess the communication data rate in each waveform. We also note that the reliability of communication depends solely on the channels and, as such, has not been considered as a performance measure in this work. In terms of implementation feasibility, the communication transmitter of all constant modulus waveforms considered in this paper can be implemented using the existing OFDM index modulation (OFDM-IM) architecture. However, the communication receiver structure depends on the particular scheme and the considered data modulation technique as explained in Section III. Therefore, we consider the computational complexity of the communication receiver to assess the feasibility of implementation for each waveform. More specifically, we use the argument of the commonly used $\mathcal{O}(\cdot)$ complexity divided by the subpulse number as the per subpulse complexity measure.
@@ III-B Waveforms with Fixed Frequency Patterns
Traditional stepped frequency waveforms, such as those with linear patterns of all available frequencies represented by the left-most time-frequency blocks in the dashed rectangle in Fig. 1, provide a deterministic control of the waveform bandwidth [8, Chapter 7]. While linear stepped frequency (LSF) waveforms suffer from high PSLs caused by large diagonal ridges in their AFs, Costas-coded stepped frequency waveforms consider well-designed frequency permutations to minimise the AF PSLs. Phase shift keying (PSK) can be independently applied to each time slot (or subpulse) of the stepped frequency waveform to embed data, which has little impact on their radar performance [8, Chapter 7]. Since a fixed frequency pattern is used, the data rate and the communications receiver are just those for pure PSK.
@@ III-C Waveforms with Variable Frequency Patterns
The data modulation can also be performed using frequency modulation, for example, using all possible frequency permutations instead of a particular one [8, Chapter 7]. An example frequency permutation is presented by the time-frequency blocks in the middle of the dashed rectangle in Fig. 1. Such a scheme maintains the deterministic control of the waveform bandwidth, while the AF PSL varies with the frequency pattern. The data rate of the permutation-based scheme depends on the total number of waveforms, which equals the factorial of the number of subpulses. The optimal communications symbol detector uses the efficient Hungarian algorithm, whose worst-case complexity for detecting all symbols in the waveform is in the order of the number of subpulses cubed. Carefully selected subsets of frequency permutations. can be considered to either improve the radar performance or decrease the communications detection error probability, though it leads to a loss of data rate. The data rate of the frequency permutation-based scheme can be boosted by incorporating subpulse-wise independent PSK with little impact on the radar performance. However, it introduces extra complexity to the optimal communications symbol detector.
To decrease this complexity, subpulse-wise independent frequency shift keying (FSK), shown by the right-most time-frequency blocks in the dashed rectangle in Fig. 1, has been considered [8, Chapter 7]. This allows the frequency to repeat, which means that deterministic control of waveform bandwidth is lost but rate and complexity are helped. Phase modulation can also be incorporated into the FSK scheme to either increase the data rate using PSK [8, Chapter 7] or decrease AF PSLs using specifically designed phase sequences [15]. For the latter, the phase sequence for each frequency sequence is obtained by solving an optimisation problem whose objective function is the PSL. Note that with the phase sequence, the independence between subpulses no longer exists. In this paper, the efficient suboptimal non-coherent detector is used for the communications receiver. This detector has the same complexity as the optimal FSK detector at the cost of a slight increase in the error rate.
Please note that most conclusions on the waveforms mentioned in Section III-B and Section III-C are drawn from [8, Chapter 7], apart from the phase optimisation one which is from [15]. For more details, please refer to these two materials.
@@ IV Performance Comparison
Since the theoretical analysis of some radar performance metrics is not available for all the signalling schemes considered, the radar performance metrics are numerically evaluated based on Monte Carlo simulations using a thousand waveform realisations for each scheme. The empirical means of all radar metrics are shown by the vertical axis values of the markers labelled in the figure legends. If the value of a particular radar metric of a scheme depends on the data, its statistics are visualised using a box plot with the edge colour the same as the corresponding marker. If two schemes have the same performance for a particular communications metric, their box plots are slightly shifted horizontally to avoid overlap, with dotted lines pointing to the same value on the horizontal axes. In each figure, we use an arrow with the text “Best” pointing to the corner indicating the most desired performance.
@@
Fig. 2 and Fig. 3 show the squared RMS bandwidth, normalised with respect to the subpulse width $T$, versus the two communications performance metrics. Although not shown here, the average run times of the communications detectors implemented via Matlab simulations follow the same order as the receiver complexities we presented in the numerical examples. The use of stepped frequency sequences, both the linear pattern and the Costas code, and frequency permutations results in stable delay local accuracy which is independent of the particular choice of transmitted waveform. The data rates and the receiver complexities for two PSK-modulated stepped frequency schemes depend only on the PSK modulation order, i.e., $4$, resulting in the best complexities and the worst rates among all discussed schemes; while the use of permutations leads to the second lowest data rate and a relatively high symbol detector complexity. Introducing PSK to the permutations significantly improves the data rate while not affecting the stable delay local accuracy, but it further increases the already high detector complexity. The FSK-only scheme and the phase-optimised FSK scheme have detector complexities much lower than the permutations-based ones and a relatively high data rate. Incorporating PSK with FSK leads to the largest data rate among all schemes and a moderate receiver complexity. Nevertheless, the subpulse-wise independent frequency modulation introduces uncontrolled delay local accuracy which is shown by the outliers of these three box plots. We observe that the FSK-based scheme with phase modulation has similar RMS bandwidth performance to that without phase modulation. This is because the initial phase terms have a negligible impact on the RMS bandwidth of an FSK-based waveform except in the rare situation where most subpulses have the same centre frequency.
@@
Figure 2: Box plots of the normalised squared RMS bandwidth vs. data rate in bits/subpulse for the seven constant modulus JCR signalling schemes. Note that we use perms. as the abbreviation of permutations in Fig. 2 - Fig. 5.
@@
Figure 3: Box plots of the normalised squared RMS bandwidth vs. the average per subpulse symbol detector complexity for the seven constant modulus JCR signalling schemes.
@@
Figure 4: Box plots of the AF PSL vs. data rate in bits/subpulse for seven constant modulus JCR signalling schemes.
Figure 5: Box plots of the AF PSL vs. the average per subpulse symbol detector complexity for seven constant modulus JCR signalling schemes.
To summarise, among all the schemes considered, if one wants the highest data rate, a relatively low communications detector complexity and moderate AF PSLs while the variance in the RMS bandwidth does not matter, one should combine FSK and PSK. A lower symbol detection complexity with similar AF PSL and RMS bandwidth performance can be achieved by using FSK only, but it reduces the data rate. Performing phase-based PSL suppression to FSK cannot stabilise the RMS bandwidth, but it leads to extremely low and stable AF PSLs and the same low symbol detector complexity and data rate as FSK only. If one wants a relatively high data rate, relatively low and stable AF PSLs and an invariant RMS bandwidth with the data modulation, one should choose PSK modulated frequency permutations, despite the fact that it has the highest symbol detector complexity. Using permutations only results in the same stable RMS bandwidth and a lower detector complexity (although it is the second highest) by slightly sacrificing the AF PSLs and dropping the data rate to a moderate level. The phase-modulated linear stepped frequencies and the Costas scheme also have stable RMS bandwidths. Note that the Costas-based scheme has the lowest AF PSL and the linear scheme has the highest, but their communications performances are similar, i.e., the lowest detector complexity but the lowest data rate.

### v2

@@ Abstract
Integrated sensing and communications (ISAC) is considered a key enabler to support application scenarios such as the Internet-of-Things (IoT) in which both communications and sensing play significant roles. Multi-carrier waveforms, such as orthogonal frequency division multiplexing (OFDM), have been considered as good candidates for ISAC due to their high communications data rate and good time bandwidth property for sensing. Nevertheless, their high peak-to-average-power-ratio (PAPR) values lead to either performance degradation or an increase in system complexity. This can make OFDM unsuitable for IoT applications with insufficient resources in terms of power, system complexity, hardware size or cost. This article provides IoT-centric constant modulus signalling scheme designs that leverage the advantage of unit PAPR and thus are more suitable in resource-limited scenarios. More specifically, several single-carrier frequency and/or phase signalling schemes are considered. A comprehensive discussion on their radar sensing and communications performance is conducted based on performance metrics including the radar ambiguity function, the bandwidth property, the data rate, and the communications receiver complexity. Results demonstrate that under the constraint of unit PAPR, the sensing-communications tardeoff can be achieved by selecting among the discussed signalling schemes. Recommendations for linking particular low-rate IoT scenarios with those signalling schemes are also provided based on the their performance.
@@ I-A Standardisation Status
Standardisation plays a crucial role in developing a consistent framework for ISAC in next generation wireless networks, within which both academia and industry can work. Recently, the industry has started laying the vision for the International Telecommunication Union Radiocommunication Sector (ITU-R) International Mobile Telecommunications-2030 (IMT-2030) towards the sixth generation of wireless communications technology (6G). ISAC has emerged as a popular theme in this vision, which led to its adoption as one of the six usage scenarios in ITU-R IMT-2030 [1]. This interest was the culmination of the progress made in researching orthogonal frequency division multiplexing (OFDM)-based radio radar sensing during the last decade, the successful introduction of radio frequency (RF) sensing in the Institute of Electrical and Electronics Engineers (IEEE) 802 standards at both low and high frequencies, and the enhancements made for the positioning framework in the Third Generation Partnership Project (3GPP). This positioning work is part of 5G New Radio (NR) to enable centimetre-level positioning as required by certain vertical industries. Following ISAC adoption in ITU-R IMT-2030, in late 2023, the European Telecommunications Standards Institute (ETSI) launched an industry specification group (ISG) on ISAC with the mission to streamline ISAC pre-standardisation research and pave the way for its specification in 6G standards. In parallel, 3GPP Technical Specification Group (TSG) on Service and System Aspects Working Group 1 (SA1) launched a feasibility study on ISAC use cases in 5G-Advanced [1]. Later in 2024, 3GPP TSG Radio Access Network Working Group 1 (RAN1) started a study item focused on extending the 3GPP baseline channel model in Technical Report (TR) 38.901 to support ISAC and hence pave the way for ISAC specification in both 5G-Advanced and 6G releases. At the recent 3GPP 6G workshop in March 2025, ISAC was confirmed as one of the key themes proposed by the industry for 3GPP to specify in the forthcoming Release 21 (2028-2029) [2]. In addition to the ongoing work of the standards development organisations, extensive pre-standardisation work for important 6G ISAC use cases has also been conducted and published by various important industrial associations. These include the 5G Automotive Association [3], 5G Alliance for Connected Industries and Automation [4] and one6G [5], who have presented various use cases in industrial and/or automated vehicle scenarios, with details of their assumptions, functional and performance requirements.
@@
In the development of the 6G framework, massive and sustainable Internet of Things (IoT) is considered as one of the key usage services for mobile users. ISAC is able to fulfil the sustainability aspect very well since it supports resource sharing between sensing and communication operations. For example, ISAC could allow transmitting a single joint signal for both sensing and communications tasks. This enables the reuse of spectrum, transmitter hardware and signal processing, which in turn reduces the size, complexity and power consumption of devices. Such efficient resource utilisation can improve the sustainability and reduce the cost of IoT networks, especially those with massive end devices.
@@ I-B Waveform Design
Various approaches have been taken to tackle this joint waveform design problem [6]. Sensing (or radar) centric methods focus on embedding information into traditional radar waveforms, including linear frequency-modulated (LFM), stepped frequency or frequency hopping (FH) multiple-input multiple-output (MIMO) waveforms. Communications-centric designs integrates the radar sensing functionality into existing communications systems. Examples include designing preambles of IEEE 802.11-based frameworks and using multi-carrier waveforms such as OFDM or orthogonal time frequency space (OTFS) for sensing. Waveforms designed from the ground up do not rely on existing radar or communications waveforms. These designs are usually formulated as optimisation problems which consider both communications and sensing performance metrics by allocating resources among temporal, spectral and/or spatial domains. It can be observed from existing methods that the time-frequency resource allocation plays a significant role in tackling the joint waveform design problem. For example, stepped frequency radar waveforms uses a single subcarrier in each time slot, while the frequencies in different time slots are different. Thus, the entire waveform occupies a large bandwidth, leading to a high range estimation resolution. Nevertheless, its multi-carrier counterparts can achieve a much higher data rate while guaranteeing the same large bandwidth. In addition, multi-carrier systems allow for trading off between communications and radar performance by adjusting the portions of subcarriers used for transmitting data and those designed for sensing. These are some of the main reasons why OFDM has been considered an important candidate for ISAC [7].
Nevertheless, many 6G systems may not have sufficient resources in terms of power, cost and size to support advanced multi-carrier techniques with the associated circuitry and components. For example, some existing systems such as Bluetooth low energy (BLE) and long-range radio (LoRA), as well as new applications introduced by 6G, such as internet of radars (IoR), smart wearables and wireless radar sensor networks (WRSNs) require low-power, inexpensive devices [8]. Therefore, single carrier modulation and low PAPR waveforms are recommended for such applications that utilise high frequency bands [9]. Given that constant envelope modulation only requires simple, low-cost circuits and operates power amplifiers at the highest efficiency point, it is ideal in low-power systems where energy efficiency and cost outweigh spectrum efficiency. As a result, in this work we focus on constant modulus waveform designs suitable for ISAC.
@@ I-C Contributions
  • We introduce multiple constant modulus frequency and/or phase signalling schemes. General discussions on the performance of these waveforms are also conducted.
  • We provide extensive numerical examples to perform a thorough comparison among the constant modulus schemes based on the introduced performance metrics. Recommendations for linking particular IoT scenarios with those signalling schemes are also provided.
Note that though we use the phrase waveform design throughout this paper, the major focus is a sub-topic of ISAC waveform design, i.e., the design of ISAC signalling or modulation schemes. More specifically, we focus on constant modulus ISAC signallings in the time-frequency domain. Therefore, existing works on other sub-topics, such as [10] for signal processing, will not be discussed in this paper.
@@ II-A Importance of Low PAPR
Nonetheless, multi-carrier waveforms that suffer from high PAPR have received significant attention as ISAC waveforms due to their high data rate. As a result, different PAPR reduction techniques such as linearisation circuits have also been proposed for ISAC systems to reduce the performance degradation [12]. However, such solutions require additional overheads of circuitry and increased complexity. This, in turn, increases the device cost or degrades the bit-error-rate (BER) performance due to OFDM signal distortion, making them unsuitable for low-cost, low-complexity and low-power scenarios. For example, existing technologies such as BLE, LoRA and IoR do not have sufficient resources in terms of power, size and cost to accommodate multi-carrier modulation with complex PAPR reduction techniques. In contrast, single-carrier constant modulus waveforms have a unit PAPR by design, without requiring extra resources to maintain the property. Hence, they can offer advantages including better performance with simplified RF components and reduced size, weight and cost, making them ideal for applications where energy efficiency and cost outweigh spectrum efficiency, e.g., low rate IoT applications with massive battery-powered devices.
@@ II-B Radar Sensing Performance
The range and velocity (or equivalently, delay and Doppler) estimation capabilities of a radar waveform can be measured in terms of local accuracy and global accuracy. Focusing on the mainlobe of the delay-Doppler profile, i.e., the ambiguity function (AF), we can analyse the local accuracy, i.e., the delay resolution and the Doppler resolution. The Doppler resolution is determined by the overall time duration of the signal and, it remains constant if the time duration of the waveforms are constant. On the other hand, the radar delay resolution is directly related to the mean square bandwidth of the radar signal under a pulse compression system. The higher the mean square bandwidth, the higher the delay resolution. Therefore, we consider the mean square bandwidth to asses the radar local performance. Focusing on the rest of the delay-Doppler plane we can analyse the global accuracy of radar sensing. It describes the accuracy of target detection in the presence of clutter and this can be measured by false alarm rate and miss-detection probability. When the transmit power remains the same, as is the case for all the waveforms discussed in this paper, the miss-detection probability will be almost the same for a given threshold. On the other hand, the false alarm rate depends on the energy distribution of the received signal. This can be evaluated using the peak sidelobe level (PSL), defined as the ratio of the magnitude of the highest sidelobe peak in the AF to the magnitude of its mainlobe peak . If the PSL is lower, it mitigates the masking effect of nearby targets at the receiver matched filter output and reduces the false alarm rate. Therefore, we consider the AF PSL to assess the radar global accuracy.
Note the aforementioned sensing metrics are useful only when the target delay and/or Doppler estimation is the sensing task. While it is supplementary for Bluetooth or LoRA, technologies such as ultra-wideband IoT and IoR require high precision ranging. Some IoR scenarios even consider conventional monostatic radar architecture for localisation, since it can be realised easily when perfect synchronisation is unavailable [8].
@@ II-C Communications Performance
From the communication perspective, the performance can be measured in terms of throughput/data rate, reliability and feasibility of implementation. Note that the communication data rate depends on the channel and can vary significantly between different communication channels. Given that the scope of this work is limited to a review of different waveforms, we consider the maximum possible data rate achievable under a perfect channel instead of focusing on data rate under different communication channels. Therefore, we consider the number of bits that can be transmitted in each sub-pulse, irrespective of the channel, to assess the communication data rate in each waveform. We also note that the reliability of communication depends solely on the channels and, as such, has not been considered as a performance measure in this work. In terms of implementation feasibility, the communication transmitter of all constant modulus waveforms considered in this paper can be implemented using the existing OFDM index modulation (OFDM-IM) architecture. However, the communication receiver structure depends on the particular scheme and the considered data modulation technique as explained in Section III. Therefore, we consider the computational complexity of the communication receiver to assess the feasibility of implementation for each modulation scheme. More specifically, we use the argument of the commonly used $\mathcal{O}(\cdot)$ complexity divided by the subpulse number as the per subpulse complexity measure.
@@ III-B Waveforms with Fixed Frequency Patterns
Traditional stepped frequency waveforms, such as those with linear patterns of all available frequencies represented by the left-most time-frequency blocks in the dashed rectangle in Fig. 1, provide a deterministic control of the waveform bandwidth [6, Section 7.1.3]. While linear stepped frequency (LSF) waveforms suffer from high PSLs caused by large diagonal ridges in their AFs, Costas-coded stepped frequency waveforms consider well-designed frequency permutations to minimise the AF PSLs. Phase shift keying (PSK) can be independently applied to each time slot (or subpulse) of the stepped frequency waveform to embed data, which has little impact on their radar performance [6, Section 7.1.3]. Since a fixed frequency pattern is used, the data rate and the communications receiver are just those for pure PSK.
@@ III-C Waveforms with Variable Frequency Patterns
The data modulation can also be performed using frequency modulation, for example, using all possible frequency permutations instead of a particular one [6, Section 7.3]. An example frequency permutation is presented by the time-frequency blocks in the middle of the dashed rectangle in Fig. 1. Such a scheme maintains the deterministic control of the waveform bandwidth, while the AF PSL varies with the frequency pattern. The data rate of the permutation-based scheme depends on the total number of waveforms, which equals the factorial of the number of subpulses. The optimal communications symbol detector uses the efficient Hungarian algorithm, whose worst-case complexity for detecting all symbols in the waveform is in the order of the number of subpulses cubed. Carefully selected subsets of frequency permutations can be considered to either improve the radar performance or decrease the communications detection error probability, though it leads to a loss of data rate [6, Section 7.4]. The data rate of the frequency permutation-based scheme can be boosted by incorporating subpulse-wise independent PSK with little impact on the radar performance [14]. However, it introduces extra complexity to the optimal communications symbol detector.
To decrease this complexity, subpulse-wise independent frequency shift keying (FSK), shown by the right-most time-frequency blocks in the dashed rectangle in Fig. 1, has been considered [15]. This allows the frequency to repeat, which means that deterministic control of waveform bandwidth is lost but rate and complexity are helped. Phase modulation can also be incorporated into the FSK scheme to either increase the data rate using PSK [6, Section 7.2] or decrease AF PSLs using specifically designed phase sequences [15]. For the latter, the phase sequence for each frequency sequence is obtained by solving an optimisation problem whose objective function is the PSL. Note that with the phase sequence, the independence between subpulses no longer exists. In this paper, the efficient suboptimal non-coherent detector is used for the communications receiver. This detector has the same complexity as the optimal FSK detector at the cost of a slight increase in the error rate.
Please note that most conclusions on the signalling schemes mentioned in Section III-B and Section III-C are drawn from [6, Chapter 7], [14] and[15]. For more details, please refer to these three materials.
@@ IV Performance Comparison
Since the theoretical analysis of some radar performance metrics is not available for all the signalling schemes considered, the radar performance metrics are numerically evaluated based on Monte Carlo simulations using $5000$ waveform realisations for each scheme. The empirical means of all radar metrics are shown by the vertical axis values of the markers labelled in the figure legends. If the value of a particular radar metric of a scheme depends on the data, its statistics are visualised using a box plot with the edge colour the same as the corresponding marker. If two schemes have the same performance for a particular communications metric, their box plots are slightly shifted horizontally to avoid overlap, with dotted lines pointing to the same value on the horizontal axes. In each figure, we use an arrow with the text “Best” pointing to the corner indicating the most desired performance.
@@
Fig. 2 and Fig. 3 show the squared RMS bandwidth, normalised with respect to the subpulse width $T$, versus the two communications performance metrics. Although not shown here, the average run times of the communications detectors implemented via Matlab simulations follow the same order as the receiver complexities we presented in the numerical examples. The use of stepped frequency sequences, both the linear pattern and the Costas code, and frequency permutations results in stable delay local accuracy which is independent of the particular choice of transmitted waveform. The data rates and the receiver complexities for two PSK-modulated stepped frequency schemes depend only on the PSK modulation order, i.e., $4$, resulting in the best complexities and the worst rates among all discussed schemes; while the use of permutations leads to the second lowest data rate and a relatively high symbol detector complexity. Introducing PSK to the permutations significantly improves the data rate while not affecting the stable delay local accuracy, but it further increases the already high detector complexity. The FSK-only scheme and the phase-optimised FSK scheme have detector complexities much lower than the permutations-based ones and a relatively high data rate. Incorporating PSK with FSK (FSK+PSK) leads to the largest data rate among all schemes and a moderate receiver complexity. Nevertheless, the subpulse-wise independent frequency modulation introduces uncontrolled delay local accuracy which is shown by the outliers of these three box plots. We observe that the FSK-based scheme with phase modulation has similar RMS bandwidth performance to that without phase modulation. This is because the initial phase terms have a negligible impact on the RMS bandwidth of an FSK-based waveform except in the rare situation where most subpulses have the same centre frequency.
@@
Figure 2: Box plots of the normalised squared RMS bandwidth vs. data rate in bits/subpulse for the seven constant modulus ISAC signalling schemes. Note that we use perms. as the abbreviation of permutations in Fig. 2 - Fig. 5.
@@
Figure 3: Box plots of the normalised squared RMS bandwidth vs. the average per subpulse symbol detector complexity for the seven constant modulus ISAC signalling schemes.
@@
Figure 4: Box plots of the AF PSL vs. data rate in bits/subpulse for seven constant modulus ISAC signalling schemes.
Figure 5: Box plots of the AF PSL vs. the average per subpulse symbol detector complexity for seven constant modulus ISAC signalling schemes.
Based on the performance evaluations in Fig. 2 to Fig. 5, Table I links the aforementioned signalling schemes with particular IoT scenarios. Since single carrier schemes could not provide as high data rate as their multi-carrier counterparts, in Table I we only consider low to medium rate scenarios, including BLE, LoRA and IoR. LoRA and BLE require low complexity communications with low (to medium) data rates [13], which can be satisfied by FSK and FSK+PSK. The uncontrolled target ranging and/or velocity estimation capability of both schemes is an insignificant issue since delay-Doppler estimation is not the main focus of LoRA and BLE. WRSNs used for target localisation prioritises high precision target sensing over communications [8] and thus, traditional radar waveforms-based schemes are recommended. In home scenarios with rich clutter (e.g., walls and furniture), clutter mitigation is prioritised over range resolution for sensing. The inter-device or device-smart gateway communications also require slightly higher data rate compared to WRSNs. Hence, the phase optimised FSK with close to ideal AF PSL and acceptable data rate is recommended. Special recommendations of using frequency permutation-based schemes are made for end devices that frequently perform uplink communications with the smart gateway. In this scenario, since the communications receiver is the gateway, slightly higher symbol detection complexity is acceptable.
TABLE I: Recommendations for linking IoT scenarios with signalling schemes.
IoT technologies Scenarios Performance requirements Recommended schemes
BLE Short range scenarios with small battery-powered devices Low to medium data rate, low complexity [13]. Target sensing is not the main focus FSK+PSK [6, Section 7.2]
LoRA Long range scenarios with massive devices Low data rate, low complexity [13]. Target sensing is not the main focus FSK [15]
IoR [8] WRSNs for target localisation High precision target sensing is prioritised. Very low data rate, low complexity Costas/LSF + PSK [6, Section 7.1.3]
Clutter-limited scenarios Clutter mitigation is prioritised for sensing. Low data rate, low complexity. FSK optimised [15]
Device to smart gateway uplink transmission Slightly higher symbol detection complexity is acceptable Frequency permutations [6, Section 7.2] Frequency permutations + PSK [14]